Can someone help me on this?
I need to add:
1/3 + 1/5 + 1/10
thanks,
Lyn

Answer by fazlerabbi(9)

1/3 + 1/5 + 1/10
= (1/3)(30/30) + (1/5)(30/30) + (1/10)(30/30)
= 10/30 + 6/30 + 3/30
= (10 + 6 + 3)/30
= 19/30
Answer is 19/30